# ADOPT: Modified Adam Can Converge with Any $\beta_{2}$ with the Optimal Rate

### v1

@@ 4 Method: Adaptive Gradient Method with the Optimal Convergence Rate
Algorithm 1 ADOPT algorithm
@@
The detailed proof and related lemmas are provided in the appendix. We also provide the convergence bound for the case of diminishing learning rate (i.e., $\alpha_{t}=\alpha/\sqrt{t}$) in the appendix, which is closer to practical situations. In that case, ADOPT also converges with the optimal rate.
@@ 5 Experiments
In the experiments, we first validate our ADOPT algorithm using a simple toy example in which Adam is known to fail to converge, and confirm our theoretical findings through numerical simulation. Secondly, we run an experiment of training a simple multi-layer perceptron (MLP) for the MNIST dataset to verify the effectiveness of our ADOPT for nonconvex optimization problems. Finally, we evaluate our ADOPT in a wide range of practical applications, including image classification, natural language processing (NLP) tasks, generative modeling, and deep reinforcement learning. Detailed experimental settings are described in the appendix.
@@ Appendix C Another Expression of ADOPT
     if $t=1$ then
        ${\bm{m}}_{t}\leftarrow{\bm{g}}_{t}/\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}$
     else
@@ Appendix E Theorems
Under Assumptions 2.1, 2.2, 2.3, and 2.5, if the objective $f$ is upper-bounded by $f_{\mathrm{sup}}$, the following holds for the ADOPT algorithm with a learning rate $\alpha_{t}=\alpha/\sqrt{t}$:
$\min_{t=1,\ldots,T}\left\{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t}\right)\right\|^{4/3}\right]^{3/2}\right\}$
$\leq\frac{3\sqrt{\max\left\{G^{2},1\right\}+\epsilon^{2}}}{2\left(\left(T+1\right)^{3/2}-1\right)}\left(\frac{f_{\mathrm{sup}}-f_{\mathrm{inf}}}{\alpha}\left(T+1\right)+\left(\frac{\sqrt{2}\alpha\beta_{1}G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)}+\frac{\alpha G^{2}L}{2\epsilon^{2}}\right)T\right)$
$\quad\ +\frac{3\sqrt{\max\left\{G^{2},1\right\}+\epsilon^{2}}}{2\left(\left(T+1\right)^{3/2}-1\right)}\left(\frac{2\sqrt{2}\beta_{1}G^{2}}{\epsilon\left(1-\beta_{1}\right)}\left(\sqrt{T+1}-1\right)+\frac{\alpha\beta_{1}^{2}G^{2}L}{\epsilon\left(1-\beta_{1}\right)^{2}}\frac{T}{T+1}\right)$
$\quad\ +\frac{3\sqrt{\max\left\{G^{2},1\right\}+\epsilon^{2}}}{2\left(\left(T+1\right)^{3/2}-1\right)}\left(\frac{2\alpha\beta_{1}^{2}G^{2}L}{\epsilon\left(1-\beta_{1}\right)^{2}}+\frac{\alpha^{2}\beta_{1}G^{2}L}{\sqrt{2}\epsilon^{2}\left(1-\beta_{1}\right)}\right)\log\left(T+1\right).$ (20)
@@ Appendix F Proofs
Proof of Theorems 4.1 and E.1.
@@
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]+\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\epsilon}-\alpha_{t}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}$ (27)
$\quad\ +\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}$
$\quad\ +\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{2\left(1-\beta_{1}\right)^{2}\epsilon^{2}}$
$\quad\ +\frac{\alpha_{t}^{2}G^{2}L}{2\epsilon^{2}}+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}}.$
@@
$\mathbb{E}\left[f\left({\bm{\phi}}_{t}\right)\right]$ $\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]-\alpha\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha^{2}G^{2}L}{2\epsilon^{2}}$ (28)
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]-\alpha\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\left(1+\beta_{1}\right)G^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}}.$ (29)
@@
$\mathbb{E}\left[f\left({\bm{\phi}}_{T}\right)\right]$ $\leq f\left({\bm{\theta}}_{0}\right)-\alpha\frac{\sum_{t=1}^{T}\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\left(1+\beta_{1}\right)G^{2}LT}{2\left(1-\beta_{1}\right)\epsilon^{2}}$ (30)
$\leq f\left({\bm{\theta}}_{0}\right)-\alpha\frac{\sum_{t=1}^{T}\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\left(1+\beta_{1}\right)G^{2}LT}{2\left(1-\beta_{1}\right)\epsilon^{2}}$ (31)
@@
$\leq\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}\left(\frac{f\left({\bm{\theta}}_{0}\right)-f_{\mathrm{inf}}}{\alpha T}+\frac{\alpha\left(1+\beta_{1}\right)G^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}}\right)$ (33)
When $\alpha_{t}=\alpha/\sqrt{t}$, the following holds for $t\geq 2$:
@@
$\mathbb{E}\left[f\left({\bm{\phi}}_{t}\right)\right]$ $\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]+\frac{\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\epsilon}-\alpha_{t}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}$
$\quad\ +\frac{\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}$
$\quad\ +\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{2\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{\alpha_{t}^{2}G^{2}L}{2\epsilon^{2}}$
$\quad\ +\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}}$ (40)
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]+\frac{\sqrt{2}\alpha\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{t^{3/2}\left(1-\beta_{1}\right)\epsilon}-\frac{\alpha}{\sqrt{t}}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}$
$\quad\ +\frac{2\alpha^{2}\beta_{1}^{2}G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}t^{2}}+\frac{\sqrt{2}\alpha^{2}\beta_{1}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}t}+\frac{\alpha^{2}\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}t^{3}}+\frac{\alpha^{2}G^{2}L}{2\epsilon^{2}t}$
$\quad\ +\frac{\alpha^{2}\beta_{1}G^{2}L}{\sqrt{2}\left(1-\beta_{1}\right)\epsilon^{2}t^{2}}$ (41)
$=\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]-\frac{\alpha}{\sqrt{t}}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}$
$\quad\ +\frac{\alpha^{2}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}}\cdot t^{-1}+\frac{\sqrt{2}\alpha\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\cdot t^{-\frac{3}{2}}$
$\quad\ +\frac{\alpha^{2}\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\sqrt{2}\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\cdot t^{-2}+\frac{\alpha^{2}\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\cdot t^{-3}.$ (42)
@@
$\frac{\sqrt{t}\ \mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}$
$\leq\frac{\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)-f\left({\bm{\phi}}_{t}\right)\right]}{\alpha}\cdot t+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{2\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\sqrt{2}\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\cdot t^{-\frac{1}{2}}$ (43)
$\quad\ +\frac{\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\sqrt{2}\left(1-\beta_{1}\right)^{2}\epsilon^{2}}t^{-1}+\frac{\alpha\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}t^{-2}$ (44)
$\sum_{t=1}^{T}\frac{\sqrt{t}\ \mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{\left(1-\beta_{2}^{T}\right)G^{2}+\epsilon^{2}}}$
$\leq\frac{f\left({\bm{\phi}}_{0}\right)-Tf\left({\bm{\phi}}_{T}\right)+\sum_{t=1}^{T-1}f\left({\bm{\phi}}_{t}\right)}{\alpha}+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}LT}{2\left(1-\beta_{1}\right)\epsilon^{2}}$
$\quad\ +\frac{\sqrt{2}\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\sum_{t=1}^{T}t^{-\frac{1}{2}}+\frac{\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\sqrt{2}\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\sum_{t=1}^{T}t^{-1}+\frac{\alpha\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\sum_{t=1}^{T}t^{-2}$ (45)
$\leq\frac{f_{\mathrm{sup}}-f_{\mathrm{inf}}}{\alpha}T+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}LT}{2\left(1-\beta_{1}\right)\epsilon^{2}}$

### v3

@@ 4 Method: Adaptive Gradient Method with the Optimal Convergence Rate
0:  Exponential decay rate $10\leq\beta_{1},\beta_{2},\beta_{3}<$ 1, small constant $\epsilon>0$
@@
Algorithm 2 Clipped ADOPT algorithm
0:  Learning rate $\left\{\alpha_{t}\right\}$, clipping value $\left\{c_{t}\right\}$, initial parameter ${\bm{\theta}}_{0}$
0:  Exponential decay rate $10\leq\beta_{1},\beta_{2}<$ 1, small constant $\epsilon>0$
  ${\bm{m}}_{0}\leftarrow{\bm{0}},{\bm{v}}_{0}\leftarrow{\bm{g}}_{0}\odot{\bm{g}}_{0}$
  for $t=1$ to $T$ do
      ${\bm{m}}_{t}\leftarrow\beta_{1}\cdot{\bm{m}}_{t-1}+\left(1-\beta_{1}\right)\mathrm{Clip}\left(\frac{{\bm{g}}_{t}}{\max\left\{\sqrt{{\bm{v}}_{t-1}},\epsilon\right\}},c_{t}\right)$
     ${\bm{\theta}}_{t}\leftarrow{\bm{\theta}}_{t-1}-\alpha_{t}{\bm{m}}_{t}$
  end for
  return $\{{\bm{\theta}}_{t}\}_{t=1}^{T}$
@@
The detailed proof and related lemmas are provided in the appendix. We also provide the convergence bound for the case of diminishing learning rate (i.e., $\alpha_{t}=\Theta(1/\sqrt{t})$) in the appendix, which is closer to practical situations. In that case, ADOPT also converges with the optimal rate.
In practice, however, the ADOPT algorithm sometimes becomes unstable especially when near-zero gradients are observed in the early phase of optimization. For example, if some elements of ${\bm{v}}_{0}$ are almost zero, the corresponding elements of ${\bm{g}}_{1}/\max\{\sqrt{{\bm{v}}_{0}},\epsilon\}$ take very large values due to the near-zero division, leading to an unstable parameter update. Such a phenomenon is typically observed when, for example, some parts of parameters (e.g., the last layer of a neural net) are initialized with zero, which is a commonly-used technique in deep learning. To avoid the near-zero division, we also propose a clipped version of ADOPT in Algorithm 2. Note that the clipping operation is applied in an element-wise manner:
$\mathrm{Clip}\left({\bm{a}},c\right)_{i}=\min\left\{\max\left\{{a}_{i},-c\right\},c\right\},$ (16)
where $c\geq 0$. Even when the clipping is incorporated, the same convergence rate can be guaranteed by properly scheduling the clipping value $c_{t}$ (see Theorems E.2 and E.3 in the appendix). Specifically, we can guarantee the convergence by setting $c_{t}=c=\Theta\left({T}^{1/4}\right)$ or $c_{t}=\Theta\left({t}^{1/4}\right)$.
@@ 5 Experiments
In the experiments, we first validate our ADOPT algorithm using a simple toy example in which Adam is known to fail to converge, and confirm our theoretical findings through numerical simulation. Secondly, we run an experiment of training a simple multi-layer perceptron (MLP) for the MNIST dataset to verify the effectiveness of our ADOPT for nonconvex optimization problems. Finally, we evaluate our ADOPT in a wide range of practical applications, including image classification, natural language processing (NLP) tasks, generative modeling, and deep reinforcement learning. Detailed experimental settings are described in the appendix. In these experiments, we use the vanilla ADOPT algorithm in Algorithm 1 and do not use the clipped version in Algorithm 2.
@@ Appendix E Theorems
Under Assumptions 2.1, 2.2, 2.3, and 2.5, if the objective $f$ is upper-bounded by $f_{\mathrm{sup}}$, the following holds for the ADOPT algorithm with a learning rate $\alpha_{t}=\Theta\left(1/\sqrt{t}\right)$:
$\min_{t=1,\ldots,T}\left\{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t}\right)\right\|^{4/3}\right]^{3/2}\right\}=\mathcal{O}\left(1/\sqrt{T}\right).$ (21)
Theorem E.2.
Under Assumptions 2.1-2.3 and 2.5, the following holds for the clipped ADOPT algorithm with a constant learning rate $\alpha_{t}=\alpha=\Theta\left(1/\sqrt{T}\right)$ when clipping value meets $c_{t}=c=\Theta\left(T^{1/4}\right)$ or $c_{t}=\Theta\left(t^{1/4}\right)$:
$\min_{t=1,\ldots,T}\left\{\mathbb{E}\left[\left\|\nabla f({\bm{\theta}}_{t-1}))\right\|^{4/3}\right]^{3/2}\right\}\leq\mathcal{O}\left(1/\sqrt{T}\right).$ (22)
Theorem E.3.
Under Assumptions 2.1, 2.2, 2.3, and 2.5, if the objective $f$ is upper-bounded by $f_{\mathrm{sup}}$, the following holds for the ADOPT algorithm with a learning rate $\alpha_{t}=\Theta\left(1/\sqrt{t}\right)$ and a clipping value $c_{t}=\Theta\left(t^{1/4}\right)$:
$\min_{t=1,\ldots,T}\left\{\mathbb{E}\left[\left\|\nabla f({\bm{\theta}}_{t-1}))\right\|^{4/3}\right]^{3/2}\right\}\leq\mathcal{O}\left(1/\sqrt{T}\right).$ (23)
@@ Appendix F Proofs
Proof of Theorems 4.1, E.2, E.1, and E.3.
We only provide the proof for the clipped ADOPT here, since the vanilla ADOPT can be seen as a special case of the clipped ADOPT with $c_{t}=\infty$.
@@
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]+\frac{\sqrt{2}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\epsilon}-\frac{\alpha_{t}}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}$ (30)
$\quad\ +\frac{2\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{2\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}$
$\quad\ +\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}$
$\quad\ +\frac{\alpha_{t}^{2}G^{2}L}{\epsilon^{2}}+\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}.$
@@
$\mathbb{E}\left[f\left({\bm{\phi}}_{t}\right)\right]$
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]-\frac{\alpha}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{2\alpha^{2}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha^{2}G^{2}L}{\epsilon^{2}}+\frac{\alpha G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (31)
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]-\frac{\alpha}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\left(1+\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha G^{4}}{2\epsilon^{3}c_{t}^{2}}.$ (32)
@@
$\mathbb{E}\left[f\left({\bm{\phi}}_{T}\right)\right]$
$\leq f\left({\bm{\theta}}_{0}\right)-\frac{\alpha}{2}\frac{\sum_{t=1}^{T}\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\left(1+\beta_{1}\right)G^{2}LT}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha G^{4}T}{2\epsilon^{3}}\sum_{t=1}^{T}c_{t}^{-2}$ (33)
$\leq f\left({\bm{\theta}}_{0}\right)-\frac{\alpha}{2}\frac{\sum_{t=1}^{T}\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}+\frac{\alpha^{2}\left(1+\beta_{1}\right)G^{2}LT}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha G^{4}T}{2\epsilon^{3}}\sum_{t=1}^{T}c_{t}^{-2}$ (34)
@@
$\leq 2\sqrt{G^{2}+\epsilon^{2}}\left(\frac{f\left({\bm{\theta}}_{0}\right)-f_{\mathrm{inf}}}{\alpha T}+\frac{\alpha\left(1+\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{G^{4}}{2\epsilon^{3}T}\sum_{t=1}^{T}c_{t}^{-2}\right).$ (36)
When $\alpha=\Theta\left(1/\sqrt{T}\right)$, the first and second terms are $\mathcal{O}\left(1/\sqrt{T}\right)$. If the clipping value is constant, and $c_{t}=c=\Theta\left(T^{1/4}\right)$, the last term is also $\mathcal{O}\left(1/\sqrt{T}\right)$:
$\frac{1}{T}\sum_{t=1}^{T}c_{t}^{-2}=\mathcal{O}\left(c^{-2}\right)=\mathcal{O}\left(1/\sqrt{T}\right).$ (37)
The same holds true when the clipping value is $c_{t}=\Theta\left(t^{1/4}\right)$, since the following holds:
$\frac{1}{T}\sum_{t=1}^{T}c_{t}^{-2}=\mathcal{O}\left(\frac{1}{T}\sum_{t=1}^{T}1/\sqrt{t}\right)=\mathcal{O}\left(1/\sqrt{T}\right).$ (38)
For the case of $\alpha_{t}=\Theta\left(1/\sqrt{t}\right)$ and $c_{t}=\Theta\left(t^{1/4}\right)$, we can set $\alpha_{t}=\alpha/\sqrt{t}$ and $c_{t}=c{t}^{1/4}$ without loss of generality. The following holds for $t\geq 2$ for this case:
@@
$\mathbb{E}\left[f\left({\bm{\phi}}_{t}\right)\right]$ $\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]+\frac{\sqrt{2}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{\left(1-\beta_{1}\right)\epsilon}-\frac{\alpha_{t}}{2}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}$
$\quad\ +\frac{2\alpha_{t-1}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}}+\frac{2\alpha_{t}\alpha_{t-1}\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}$
$\quad\ +\frac{\left(\alpha_{t-1}-\alpha_{t}\right)^{2}\beta_{1}^{2}\left(1-\beta_{1}^{t-1}\right)G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}+\frac{\alpha_{t}^{2}G^{2}L}{\epsilon^{2}}$
$\quad\ +\frac{\alpha_{t}\left(\alpha_{t-1}-\alpha_{t}\right)\beta_{1}\sqrt{1-\beta_{1}^{t-1}}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha_{t}G^{4}}{2\epsilon^{3}c_{t}^{2}}$ (44)
$\leq\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]+\frac{{2}\alpha\beta_{1}\left(1-\beta_{1}^{t-1}\right)G^{2}}{t^{3/2}\left(1-\beta_{1}\right)\epsilon}-\frac{\alpha}{2\sqrt{t}}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}$
$\quad\ +\frac{4\alpha^{2}\beta_{1}^{2}G^{2}L}{\epsilon^{2}\left(1-\beta_{1}\right)^{2}t^{2}}+\frac{2\sqrt{2}\alpha^{2}\beta_{1}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}t}+\frac{2\alpha^{2}\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}t^{3}}+\frac{\alpha^{2}G^{2}L}{\epsilon^{2}t}$
$\quad\ +\frac{\sqrt{2}\alpha^{2}\beta_{1}G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}t^{2}}+\frac{\alpha G^{4}}{2c^{2}\epsilon^{3}{t}}$ (45)
$=\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)\right]-\frac{\alpha}{2\sqrt{t}}\frac{\mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|_{i}^{4/3}\right]^{3/2}}{\sqrt{G^{2}+\epsilon^{2}}}$
$\quad\ +\left(\frac{\alpha^{2}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{\alpha G^{4}}{2c^{2}\epsilon^{3}}\right)\cdot t^{-1}+\frac{{2}\alpha\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\cdot t^{-\frac{3}{2}}$
$\quad\ +\frac{\sqrt{2}\alpha^{2}\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\cdot t^{-2}+\frac{2\alpha^{2}\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\cdot t^{-3}.$ (46)
@@
$\frac{\sqrt{t}\ \mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{2\sqrt{G^{2}+\epsilon^{2}}}$
$\leq\frac{\mathbb{E}\left[f\left({\bm{\phi}}_{t-1}\right)-f\left({\bm{\phi}}_{t}\right)\right]}{\alpha}\cdot t+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{G^{4}}{2c^{2}\epsilon^{3}}+\frac{{2}\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\cdot t^{-\frac{1}{2}}$ (47)
$\quad\ +\frac{\sqrt{2}\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}t^{-1}+\frac{2\alpha\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}t^{-2}.$ (48)
Telescoping it for $t=1,\ldots,T$, we have
$\sum_{t=1}^{T}\frac{\sqrt{t}\ \mathbb{E}\left[\left\|\nabla f\left({\bm{\theta}}_{t-1}\right)\right\|^{4/3}\right]^{3/2}}{2\sqrt{G^{2}+\epsilon^{2}}}$
$\leq\frac{f\left({\bm{\phi}}_{0}\right)-Tf\left({\bm{\phi}}_{T}\right)+\sum_{t=1}^{T-1}f\left({\bm{\phi}}_{t}\right)}{\alpha}+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}LT}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{G^{4}T}{2c^{2}\epsilon^{3}}$
$\quad\ +\frac{{2}\beta_{1}G^{2}}{\left(1-\beta_{1}\right)\epsilon}\sum_{t=1}^{T}t^{-\frac{1}{2}}+\frac{\sqrt{2}\alpha\beta_{1}\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\sum_{t=1}^{T}t^{-1}+\frac{2\alpha\beta_{1}^{2}G^{2}L}{\left(1-\beta_{1}\right)^{2}\epsilon^{2}}\sum_{t=1}^{T}t^{-2}$ (49)
$\leq\frac{f_{\mathrm{sup}}-f_{\mathrm{inf}}}{\alpha}T+\frac{\alpha\left(1+\left(2\sqrt{2}-1\right)\beta_{1}\right)G^{2}LT}{\left(1-\beta_{1}\right)\epsilon^{2}}+\frac{G^{4}T}{2c^{2}\epsilon^{3}}$